# Media Coverage of War Victims: Journalistic Biases in Reporting on Israel and Gaza

### v2

@@ Results
The imbalance becomes even clearer when looking at the distribution of child-related stories within each outlet (Supplementary Table 5b). If each child’s life were given equal weight, nearly all individualized stories about children (over 99%) would focus on Palestinian victims. Instead, Western outlets devoted a strikingly large share of their child coverage to Israeli children: BBC 44%, CNN 43%, and NYT 51%. This pattern reflects a form of false balance: by the end of the first year of war, 38 Israeli [57] and 14,000 Palestinian children [7] had been killed, a disparity likely even greater for other hardships such as injuries, displacement, and famine. If casualty numbers alone were used as a baseline, individualized stories on Israeli children would be expected to account for less than 1% of all child references. AJE came closest to reflecting the underlying disparity, with 91% of its child stories about Palestinians. The most extreme deviation is observed in the New York Times, where Israeli and Palestinian children were given nearly equal attention, despite the enormous difference in the actual number of victims.
Quantifying the Human Cost of War: the Importance of Numbers
@@
The reporting of victim numbers in news articles is crucial for conveying the magnitude of casualties and the impact of violence on communities. Quantitative data not only provides readers with a clear understanding of the scale of incidents but also shapes public perception and policy responses to crime and victimization. Research indicates that the inclusion of victim statistics can significantly influence how audiences interpret the severity of an issue, thereby fostering a sense of urgency and empathy towards affected populations [58]. For instance, when media outlets report on violent crimes, the number of victims can serve as a stark reminder of the human cost associated with such events, prompting discussions on prevention and intervention strategies [59]. Furthermore, the reporting of victim numbers can also reflect broader societal attitudes towards certain groups, as disparities in reporting are likely to lead to the devaluation of victims from marginalized communities [60]. This selective emphasis on victim statistics not only affects public discourse but can also influence legislative and law enforcement priorities, underscoring the media’s role in shaping narratives around crime and victimhood [61]. Thus, accurate and comprehensive reporting of victim numbers is essential for fostering informed public dialogue and effective policy-making.
@@
Recognizing the importance of reporting numbers in conflict scenarios, we employed an LLM to analyze all articles and extract instances containing numerical references to civilians being harmed in any capacity, such as “100 civilians were killed in that strike.” For clarity and convenience, these instances will be referred to as Civilian Victim Numbers (CVNs). A detailed explanation of how this information was extracted and validated can be found in Supplementary Note 3. We begin by examining the total number of articles containing at least one CVN. Figure 3a presents a bar plot displaying the number of articles for each of the four media sources. Each source is represented by three bars: one showing the total number of articles (grey), one for articles mentioning CVNs related to the Palestinian side (green) and one for the Israeli side (blue). Notable differences in reporting styles emerge when comparing AJE with the Western news outlets. AJE reports twice as many articles with Palestinian CVNs than Israeli, while BBC and CNN report on both sides almost equally. Furthermore, NYT presents an interesting case: although the absolute number of its articles mentioning Palestinian CVNs (2,705) is comparable to AJE’s (2,847), the proportion is markedly lower (46% for NYT versus 74% for AJE). At the same time, NYT reports more on Israeli CVNs in absolute terms than any other outlet, whereas the BBC devotes the highest proportion of its coverage to Israeli CVNs (59%).
@@
Given the significant disparity in civilian casualties reported by October 7, 2024 (over 41,689 Palestinians versus 1,200 Israelis killed, and 96,625 Palestinians versus 5,432 Israelis injured, as documented by the United Nations [4]), this imbalance should logically be reflected in media reporting, with greater attention directed toward the side experiencing more severe and prolonged impacts. To evaluate this, we assessed the proportion of CVNs specifically related to casualties (for a breakdown of the different types of CVNs, refer to Supplementary Table 6). Our findings confirm that the majority of CVNs were indeed linked to casualty-related statistics. Moreover, we observe that, across all outlets, Palestinian CVNs outnumber Israeli CVNs. However, in Western media specifically, Israeli CVNs still amount to more than half the number of Palestinian CVNs, despite the vast disparity in actual casualties. Once again, this points to a form of false balance, where Israeli suffering is elevated to a level of prominence in coverage that does not reflect the underlying reality.
@@
False Balance in Coverage: A Baseline Comparison of Casualty-Related Numbers
@@
The comparison between the expected number of casualty-related CVN mentions, based on the baseline model, and the actual number of casualty-related CVN mentions for each side is visualized over time for each news source in Supplementary Figure 5. A consistent trend emerges across all outlets: Israeli CVNs appear to be overreported, while Palestinian CVNs seem to be underreported. However, comparing the magnitude of these gaps, between expected and actual mentions, across different news sources is challenging. Just how much overreporting or underreporting, and thus potential bias, exists relative to the baseline? To address this, we calculate the weekly proportion of expected mentions allocated to each side based on the baseline model for each news outlet.
@@
Building on the previous observations, we conducted an analysis to investigate the possibility of selective citation of sources for CVNs. We utilized an LLM to evaluate each sentence containing a CVN, prompting it to determine whether a source was cited (for details, see Supplementary Note 4). Subsequently, we examined the frequency of cited sources for CVNs on each side and compared patterns across different news outlets (examples of the most popular ways to cite sources for each side can be found in Supplementary Tables 7 and 8). The results, detailed in Supplementary Table 9, highlight distinct differences in reporting practices. AJE demonstrated a balanced approach, citing sources with almost equal frequency for both Palestinian and Israeli CVNs, indicating a consistent standard for source attribution. In contrast, Western news outlets were significantly more prone to cite sources when reporting on Palestinian CVNs than Israeli ones. U.S.-based outlets, in particular, referenced sources for Palestinian data nearly twice as often as for Israeli figures. BBC, on the other hand, reported Israeli CVNs with the least source attribution, at only 16.9% of instances, while attributing sources to Palestinian CVNs three times more frequently. These disparities raise questions about potential biases in validating casualty figures based on the context.
Creating Skepticism
@@
Next, we analyzed how referenced CVNs were phrased to determine whether they were presented in a “casting doubt” style (details can be found in Supplementary note 5). Using an LLM, we examined full sentences containing CVNs for phrases implying skepticism or questioning the source or the reported numbers. These doubt-casting phrases were categorized into two main types: (1) Source Doubting, i.e. phrases that question the credibility of the reporting source, thereby indirectly doubting the reported numbers, such as: “Israel’s relentless airstrikes have killed more than 8,000 people, according to the Hamas-run health ministry[74] and “The health ministry in Gaza, which is controlled by Hamas, estimates that 11,000 civilians have been killed there over the last month,” [75] and (2) Uncertainty in Numbers, i.e. phrases that cast doubt specifically on the reported numbers, such as: “…dozens of civilians were reportedly killed when jets struck buildings…” [76]. For a complete list of doubt-casting phrases and their categories, see Supplementary Table 10. Notably, our analysis found no source-doubting phrases on Israeli-reported numbers, even after using a second independent LLM prompt designed specifically to detect such instances. This suggests that doubt-casting techniques are disproportionately applied to Palestinian-reported numbers. While this study does not aim to verify the accuracy of these numbers, its purpose is to evaluate, through numerical analysis, whether both sides are presented equally in news coverage. We explore this further below.
@@
After identifying doubt-casting phrases, we analyzed their occurrences for each side (Palestine vs. Israel) across four news sources. The results, summarized in Figure 3c, reveal that BBC overwhelmingly leads in using doubt-casting phrases, with 915 CVN mentions, 98% of which target Palestinians, averaging over two CVNs doubted daily during the first year. This far exceeds CNN (145), NYT (131), and AJE (72). Common Source Doubting phrases like “Hamas-run health ministry” were heavily used by BBC, CNN, and NYT, whereas analogous expressions for Israel (such as “Knesset-run health ministry”) were absent. CNN and NYT also favored phrases like “Hamas-controlled enclave.” For Uncertainty in Numbers, terms such as “reportedly” and “claim” were predominantly used to describe Palestinian casualties by Western outlets, whereas AJE applied them more evenly across both sides (see Supplementary Tables 11–14 for detailed lists of phrases by outlet). Western sources attributed 71–90% of their cast-doubting phrases to doubting the sources of Palestinian CVNs, compared to 21% for AJE. Additionally, AJE was the only outlet to use Uncertainty in Numbers phrases in a balanced manner: 53% referred to Palestinians and 26% to Israelis, a proportion consistent with the fact that AJE published roughly twice as many articles about Palestinian CVNs as about Israeli ones. In contrast, the other outlets applied such language almost exclusively to Palestinians, with 3% or less referring to Israelis. These findings underscore in particular BBC’s significant reliance on doubt-casting phrases, particularly when reporting on Palestinian casualties, compared to other sources.
@@
To conclude this section of our analysis, we examined how frequently media outlets referenced children in their reporting of civilian victim numbers (CVNs) for both Palestinians and Israelis (Supplementary Figures 6 and 7, and Supplementary Table 15). To contextualize these proportions, we benchmark them in Supplementary Table 15 against the approximate share of children among civilian deaths on each side: around 5% for Israeli civilians on October 7 (and roughly 3% when including soldiers, based on public compilation) [80, 57], compared to 40–44% for Palestinian civilians over the first year of the war (October 7, 2023–October 7, 2024) [81, 82].
@@ Discussion
First, unlike AJE, Western media (BBC, CNN, and NYT) tend to emphasize the individuality of Israeli victims experiencing hardship, while treating Palestinian victims more frequently as a collective. This is evident in the higher ratio of individualized to group-level reporting for the Israeli compared to the Palestinian side. The only outlet showing a more balanced pattern is AJE; here the prevalence of individualized (vs. group-level) narratives is very similar for the two sides. These findings suggest that even highly professional (and relatively “liberal”) Western media tend to downplay the individuality of Palestinian victims experiencing the traumatic (and often fatal) consequences of war, while acknowledging, to a greater extent, the personal suffering of Israeli victims. Based on previous literature on the identifiable victim effect [21], the Western reporting pattern is likely to elicit greater empathy and compassion for Israeli than for Palestinian victims. Although testing the impact of this reporting bias on the readership was beyond the aims of the present study, this appears to be a worthwhile endeavor for future research.
@@
The second systematic bias in Western media (and to a lesser degree in AJE) was the tendency to equate Israeli e Palestinian suffering despite the stark difference in all forms of hardship (casualties, displacement, arbitrary detention and kidnapping, destruction of the health, school, and university systems, etc.). In the absence of bias, news reports should reflect the actual gaps in human suffering. To give only one example, if Israeli child deaths are less than 1% and Palestinian children more than 99% of the total number of children killed during the first year of war, news reports would roughly reflect this discrepancy. In contrast to this assumption, we find a systematic deviation from actual discrepancies that emerges consistently across observations: (a) In Western media, 47% of individual narratives on children referred to Israeli children who constitutes less than 1% of the victims; (b) Including all age groups, the ratio of Palestinian to Israeli casualty-related individual stories was approximately 1:1 in Western media, despite an actual casualty ratio of 20:1. (c) Especially in Western media, Israeli Civilian Victim Numbers are over-reported and Palestinian Numbers under-reported compared to what would be expected based on the actual number of victims and the distribution over time. Together, these reporting patterns insinuate a false premise of symmetry, where, in reality, the entity and prevalence of human suffering is blatantly asymmetrical.
@@
How did BBC, CNN, and The NYT reporting convey an impression of (false) balance despite the extreme disparities in military power and human suffering? One contributing factor was the frequent return to the events of October 7th, even months later and in the absence of new incidents involving Israeli victims. Our analysis suggests that reiterations of October 7 were particularly pronounced during periods when significant events occurred on both sides, often overshadowing coverage of Palestinian suffering and diverting attention from the toll in Gaza. This practice, which contradicts the common “agenda decay” phenomenon, not only creates an illusion of balance, but also reiterates the idea that the 7th of October was the starting point and “cause” of the subsequent events. Importantly, we do not intend to suggest that the reiteration of October 7th was part of a deliberate strategy pursued by the analyzed media, however, even such unintentional biases may have contributed to the neglect of Palestians’ victimhood in public perception. Biases can arise from common and often automatic cognitive processes. As previously discussed, a single mass-casualty event tends to attract disproportionate attention and to be remembered more vividly than a higher but gradually accumulating death toll. In addition, events that occur first in time or space typically receive greater attention than those that follow. Together, these mechanisms may partly explain why the October 7th massacre remained a persistent reference point in subsequent reporting. At the same time, perfect parity in coverage should not be expected, given the profound asymmetry in casualties between the two sides and the unequal number of conflict-related events that could be reported.
@@
This reporting pattern resembles what is generally referred to as “false balance” or “false equivalence”. False balance occurs when equal space or attention (e.g. airtime) is given to two sides that, in reality, differ in validity, such as climate skeptics and climate scientists [83] or pro- and anti-vaccination proponents [84, 71]. Extending this concept to war journalism, false balance can manifest when unequal parties in a conflict are portrayed as equally culpable or their actions as equally justified. Regarding victims, a false balanced perspective may portray different sides involved in a conflict as equally affected by the conflict. This can obscure asymmetries in power, suffering, and in violations of international law, leading audiences to perceive the conflict as an intractable clash of opposing but equivalent sides. As one of the main consequences of false balance is the production of uncertainty [85], it is plausible that, in the context of war journalism, falsely balanced portrayals reduce public recognition of victimization or aggression, thereby dampening calls for accountability and action. Whether this applies to the present case remains an open question for future research.
@@
A third finding concerns the presumed reliability of the reported information and the trustworthiness of the source providing it. Unlike AJE, Western journalists were found to engage in strategies aimed at planting seeds of doubt in their readership when reporting on Palestinian victims. When quantifying the human cost of war for civilian victims, Western journalists tended to cite sources more frequently for Palestinian than for Israeli victims, as if Israeli (but not Palestinian) statistics can be taken for granted. This interpretation is further corroborated by a systematic and asymmetrical “casting doubt” style that questions the veridicality of data predominantly when referring to Palestinian victims (such as “reportedly”, “claims”, “according to the Hamas-run health ministry”, and the like). By discrediting sources in a subtle way, journalists undermine the reader’s trust in the information provided and in the reputation of the source.
@@
Together, the present media analysis has identified multiple biases in Western media that, together, may have contributed to the global divide in public opinion and to the delay in public outrage in the Western world in front of the unfolding genocide. Somewhat surprisingly, we did not discover a complementary pro-Palestinian bias in AJE, which - with a few exceptions, like the underreporting of Israeli children CVNs - displayed a relatively balanced reporting style.

### v3

@@ Results
The imbalance becomes even clearer when looking at the distribution of child-related stories within each outlet (Supplementary Table 5b). If each child’s life were given equal weight, nearly all individualized stories about children (over 99%) would focus on Palestinian victims. Instead, Western outlets devoted a strikingly large share of their child coverage to Israeli children: BBC 44%, CNN 43%, and NYT 51%. This pattern reflects a form of equalization bias: by the end of the first year of war, 38 Israeli [64] and 14,000 Palestinian children [8] had been killed, a disparity likely even greater for other hardships such as injuries, displacement, and famine. If casualty numbers alone were used as a baseline, individualized stories on Israeli children would be expected to account for less than 1% of all child references. AJE came closest to reflecting the underlying disparity, with 91% of its child stories about Palestinians. The most extreme deviation is observed in the New York Times, where Israeli and Palestinian children were given nearly equal attention, despite the enormous difference in the actual number of victims.
Quantitative Reporting as an Indicator of Equalizing Bias and One-Sided Doubt-Casting
Quantitative reporting provides a second lens through which media bias may manifest. Whereas narrative representation concerns whose suffering is made visible through individualized storytelling, quantitative reporting examines how numerical information is allocated, sourced, and qualified across the two sides. This dimension allows us to assess two forms of bias: (i) equalizing bias, reflected in the distribution of casualty-related numbers that do not mirror the underlying asymmetry in victimization, and (ii) one-sided doubt-casting, reflected in selective source attribution or linguistic uncertainty applied disproportionately to one side’s figures. Having analyzed biases in narrative representation, we next assess how they manifest in the numerical reporting of the four news outlets.
@@
Because numerical information plays a central role in shaping how audiences understand the scale and distribution of harm, disparities in how numbers are presented offer an important window into these biases. Reporting of victim numbers in news articles is crucial for conveying the magnitude of casualties and the impact of violence on communities. Quantitative data not only provides readers with a clear understanding of the scale of incidents but also shapes public perception and policy responses to crime and victimization. Research indicates that the inclusion of victim statistics can significantly influence how audiences interpret the severity of an issue, thereby fostering a sense of urgency and empathy towards affected populations [65]. For instance, when media outlets report on violent crimes, the number of victims can serve as a stark reminder of the human cost associated with such events, prompting discussions on prevention and intervention strategies [66]. Furthermore, the reporting of victim numbers can also reflect broader societal attitudes towards certain groups, as disparities in reporting are likely to lead to the devaluation of victims from marginalized communities [67]. This selective emphasis on victim statistics not only affects public discourse but can also influence legislative and law enforcement priorities, underscoring the media’s role in shaping narratives around crime and victimhood [68]. Thus, accurate and comprehensive reporting of victim numbers is essential for fostering informed public dialogue and effective policy-making.
Recognizing the importance of reporting numbers in conflict scenarios, we employed an LLM to analyze all articles and extract instances containing numerical references to civilians being harmed in any capacity, such as “100 civilians were killed in that strike.” For clarity and convenience, these instances will be referred to as Civilian Victim Numbers (CVNs). A detailed explanation of how this information was extracted and validated can be found in Supplementary Note 4. These extracted CVNs allow us to examine whether numerical attention is allocated proportionally across the two sides (equalizing bias) and whether numerical claims are treated differently in terms of source attribution or skepticism (doubt-casting bias). We begin by examining the total number of articles containing at least one CVN.
@@
Figure 3a presents a bar plot displaying the number of articles for each of the four media sources. Each source is represented by three bars: one showing the total number of articles (grey), one for articles mentioning CVNs related to the Palestinian side (green) and one for the Israeli side (blue). Notable differences in reporting styles emerge when comparing AJE with the Western news outlets. AJE reports twice as many articles with Palestinian CVNs than Israeli, while BBC and CNN report on both sides almost equally. Furthermore, NYT presents an interesting case: although the absolute number of its articles mentioning Palestinian CVNs (2,705) is comparable to AJE’s (2,847), the proportion is markedly lower (46% for NYT versus 74% for AJE). At the same time, NYT reports more on Israeli CVNs in absolute terms than any other outlet, whereas the BBC devotes the highest proportion of its coverage to Israeli CVNs (59%).
@@
Given the significant disparity in civilian casualties reported by October 7, 2024 (over 41,689 Palestinians versus 1,200 Israelis killed, and 96,625 Palestinians versus 5,432 Israelis injured, as documented by the United Nations [4]), this imbalance should logically be reflected in media reporting, with greater attention directed toward the side experiencing more severe and prolonged impacts. To evaluate this, we assessed the proportion of CVNs specifically related to casualties (for a breakdown of the different types of CVNs, refer to Supplementary Table 6). Our findings confirm that the majority of CVNs were indeed linked to casualty-related statistics. Moreover, we observe that, across all outlets, Palestinian CVNs outnumber Israeli CVNs. However, in Western media specifically, Israeli CVNs still amount to more than half the number of Palestinian CVNs, despite the vast disparity in actual casualties. Once again, this points to a form of equalizing bias, where Israeli suffering is elevated to a level of prominence in coverage that does not reflect the underlying reality.
@@
Equalizing Bias in Coverage: Baseline Comparison of Casualty-Related Numbers
@@
By comparing expected versus observed numerical attention, the model allows us to quantify the degree of equalizing bias for each outlet. The comparison between the expected number of casualty-related CVN mentions, based on the baseline model, and the actual number of casualty-related CVN mentions for each side is visualized over time for each news source in Supplementary Figure 7. A consistent trend emerges across all outlets: Israeli CVNs appear to be overreported, while Palestinian CVNs seem to be underreported. However, comparing the magnitude of these gaps between expected and actual mentions across different news sources is challenging. Just how much overreporting or underreporting—and thus potential bias—exists relative to the baseline? To address this, we calculate the weekly proportion of expected mentions allocated to each side based on the baseline model for each news outlet.
@@
Building on the previous observations, we conducted an analysis to investigate the possibility of selective citation of sources for CVNs. We utilized an LLM to evaluate each sentence containing a CVN, prompting it to determine whether a source was cited (for details, see Supplementary Note 5). Subsequently, we examined the frequency of cited sources for CVNs on each side and compared patterns across different news outlets (examples of the most popular ways to cite sources for each side can be found in Supplementary Tables 7 and 8). The results, detailed in Supplementary Table 9, highlight distinct differences in reporting practices. AJE demonstrated a balanced approach, citing sources with almost equal frequency for both Palestinian and Israeli CVNs, indicating a consistent standard for source attribution. In contrast, Western news outlets were significantly more prone to cite sources when reporting on Palestinian CVNs than Israeli ones. U.S.-based outlets, in particular, referenced sources for Palestinian data nearly twice as often as for Israeli figures. BBC, on the other hand, reported Israeli CVNs with the least source attribution, at only 16.9% of instances, while attributing sources to Palestinian CVNs three times more frequently. These disparities raise questions about potential biases in validating casualty figures based on the context.
Creating Skepticism through Doubt-Casting
@@
Next, we analyzed how referenced CVNs were phrased to determine whether they were presented in a “casting doubt” style (details can be found in Supplementary Note 6). Using an LLM, we examined full sentences containing CVNs for phrases implying skepticism or questioning the source or the reported numbers. These doubt-casting phrases were categorized into two main types: (1) Source Doubting, i.e. phrases that question the credibility of the reporting source, thereby indirectly doubting the reported numbers, such as: “Israel’s relentless airstrikes have killed more than 8,000 people, according to the Hamas-run health ministry[80] and “The health ministry in Gaza, which is controlled by Hamas, estimates that 11,000 civilians have been killed there over the last month,” [81] and (2) Uncertainty in Numbers, i.e. phrases that cast doubt specifically on the reported numbers, such as: “…dozens of civilians were reportedly killed when jets struck buildings…” [82]. For a complete list of doubt-casting phrases and their categories, see Supplementary Table 10. Notably, our analysis found no source-doubting phrases on Israeli-reported numbers, even after using a second independent LLM prompt designed specifically to detect such instances. This suggests that doubt-casting techniques are disproportionately applied to Palestinian-reported numbers. While this study does not aim to verify the accuracy of these numbers, its purpose is to evaluate, through numerical analysis, whether both sides are presented equally in news coverage. We explore this further below.
@@
After identifying doubt-casting phrases, we analyzed their occurrences for each side (Palestine vs. Israel) across four news sources. The results, summarized in Figure 3c, reveal that BBC overwhelmingly leads in using doubt-casting phrases, with 951 CVN mentions, 98% of which target Palestinians, averaging over two CVNs doubted daily during the first year. This far exceeds CNN (176), NYT (166), and AJE (93). Common Source Doubting phrases like “Hamas-run health ministry” were heavily used by BBC, CNN, and NYT, whereas analogous expressions for Israel (such as “Knesset-run health ministry”) were absent. CNN and NYT also favored phrases like “Hamas-controlled Gaza.” For Uncertainty in Numbers, terms such as “reportedly” , “allegedly”, and “claim” were predominantly used to describe Palestinian casualties by Western outlets, whereas AJE applied them more evenly across both sides (see Supplementary Tables 11-14 for detailed lists of phrases by outlet). Western sources attributed 77-89% of their cast-doubting phrases to doubting the sources of Palestinian CVNs, compared to 28% for AJE. Additionally, AJE was the only outlet to use Uncertainty in Numbers phrases in a balanced manner: 49% referred to Palestinians and 23% to Israelis, a proportion consistent with the fact that AJE published roughly twice as many articles about Palestinian CVNs as about Israeli ones. In contrast, the other outlets applied such language almost exclusively to Palestinians, with 2% or less referring to Israelis. These findings underscore in particular BBC’s significant reliance on doubt-casting phrases, particularly when reporting on Palestinian casualties, compared to other sources.
@@
To conclude this section of our analysis, we examined how frequently media outlets referenced children in their reporting of civilian victim numbers (CVNs) for both Palestinians and Israelis (Supplementary Figures 8 and 9, and Supplementary Table 15). To contextualize these proportions, we benchmark them in Supplementary Table 15 against the approximate share of children among civilian deaths on each side: around 5% for Israeli civilians on October 7 (and roughly 3% when including soldiers, based on public compilation) [32, 64], compared to 40–44% for Palestinian civilians over the first year of the war (October 7, 2023–October 7, 2024) [86, 87].
@@ Discussion
First, unlike AJE, Western media (BBC, CNN, and NYT) tend to emphasize the individuality of Israeli victims experiencing hardship, while treating Palestinian victims more frequently as a collective. This is evident in the higher ratio of individualized to group-level reporting for the Israeli compared to the Palestinian side. The only outlet showing a more balanced pattern is AJE; here the prevalence of individualized (vs. group-level) narratives is very similar for the two sides. These findings suggest that even highly professional (and relatively “liberal”) Western media tend to downplay the individuality of Palestinian victims experiencing the traumatic (and often fatal) consequences of war, while acknowledging, to a greater extent, the personal suffering of Israeli victims. This finding confirms what Grimm et al. (2025) [27] observed in a small-scale qualitative discourse analysis of 40 articles published in four German newspapers during the first 7 weeks of war, namely that Israeli victims tended to receive more personalized coverage than Palestinian victims. Based on previous literature on the identifiable victim effect [22], the Western reporting pattern is likely to elicit greater empathy and compassion for Israeli than for Palestinian victims. Although testing the impact of this reporting bias on the readership was beyond the aims of the present study, this appears to be a worthwhile endeavor for future research.
@@
The second systematic bias in Western media (and to a lesser degree in AJE) was the tendency to equate Israeli to Palestinian suffering despite the stark difference in all forms of hardship (casualties, displacement, arbitrary detention and kidnapping, destruction of the health, school, and university systems, etc.), We termed this pattern “equalizing bias”. In the absence of bias, news reports should reflect the actual gaps in human suffering. To give only one example, if Israeli child deaths are less than 1% and Palestinian children more than 99% of the total number of children killed during the first year of war, news reports would roughly reflect this discrepancy. In contrast to this assumption, we find a systematic deviation from actual discrepancies that emerges consistently across observations: (a) In Western media, 47% of individual narratives on children referred to Israeli children who constitutes less than 1% of the victims; (b) Including all age groups, the ratio of Palestinian to Israeli casualty-related individual stories was approximately 1:1 in Western media, despite an actual casualty ratio of 20:1; (c) Especially in Western media, Israeli Civilian Victim Numbers are over-reported and Palestinian Numbers under-reported compared to what would be expected based on the actual number of victims and the distribution over time. Together, these reporting patterns insinuate a false premise of symmetry, where, in reality, the entity and prevalence of human suffering is blatantly asymmetrical. This raises the question of whether the equalizing bias is unique to the Israeli-Palestinian conflict or whether it constitutes a more general bias likely to occur whenever journalists (and possibly laypeople) identify with the dominant side of an asymmetrical conflict. What argues in favor of the latter interpretation is our experimental study replicating the equalizing bias among lay people acting as journalists in an imaginary, greatly tilted conflict.
@@
How did BBC, CNN, and the NYT reporting convey an impression of (false) balance despite the extreme disparities in military power and human suffering? One contributing factor was the frequent return to the events of October 7th, even months later and in the absence of new incidents involving Israeli victims. Our analysis suggests that reiterations of October 7 were particularly pronounced during periods when significant events occurred on both sides, often overshadowing coverage of Palestinian suffering and diverting attention from the toll in Gaza. This practice, which contradicts the common “agenda decay” phenomenon, not only creates an illusion of balance, but also reiterates the idea that the 7th of October was the starting point and “cause” of the subsequent events. Importantly, we do not intend to suggest that the reiteration of October 7th was part of a deliberate strategy pursued by the analyzed media, however, even such unintentional biases may have contributed to the neglect of Palestians’ victimhood in public perception. Biases can arise from common and often automatic cognitive processes. As previously discussed, a single mass-casualty event tends to attract disproportionate attention and to be remembered more vividly than a higher but gradually accumulating death toll. In addition, events that occur first in time or space typically receive greater attention than those that follow. Together, these mechanisms may partly explain why the October 7th massacre remained a persistent reference point in subsequent reporting. At the same time, perfect parity in coverage should not be expected, given the profound asymmetry in casualties between the two sides and the unequal number of conflict-related events that could be reported. Yet these cognitive mechanisms alone can hardly explain the almost perfect parity in coverage we observe, particularly in light of the profound asymmetry in casualties and the greatly unequal number of conflict-related events that could be reported.
@@
We interpret the reporting pattern observed in Western media as an instance of what we have termed the equalizing bias, namely the tendency to portray two groups as equally affected by a conflict despite stark differences in suffering. Although this is beyond the scope of the present media analysis, the equalizing of unequal parties may not only obscure asymmetries in suffering, but also in power and in culpability (such as violations of international law), leading audiences to (mis-)perceive the conflict as an intractable clash of opposing but equivalent sides. As one of the main consequences of false equivalence is the production of uncertainty [88], it is plausible that, in the context of war journalism, equalizing portrayals will reduce public recognition of victimization and aggression, thereby dampening calls for accountability and action. Whether the equalizing bias does, indeed, have these effects remains an open question for future research.
Furthermore, why would Western journalists portray Israeli suffering at par with Palestinian suffering? We can envisage at least three, not mutually exclusive reasons: First, the equalization bias may reflect the journalistic norm of pursuing balance and impartiality (although the same desire for balance was not visible in the choice of individuating and doubt-casting reporting). This interpretation, however, was not supported by our experimental simulation study in which the endorsement of journalistic norms failed to predict the magnitude of the equalizing bias. Second, to the degree that Western reporters may identify more with the Israeli than the Palestinian side [28], they may engage in what is known as “competitive victimhood”, namely the tendency to claim that one’s own group has suffered more than a relevant outgroup [89]. In particular, competitive victimhood provides a tool for regaining moral ground whenever a group is accused of inflicting harm to the outgroup (Sullivan et al., 2012 [90]). This interpretation is supported by our experimental findings, which show that identification with the stronger country significantly predicted the equalizing bias and fully mediated the effect of experimental condition on victim selection, suggesting that group identification alone is sufficient to generate competitive victimhood dynamics. The third and most disquieting possibility is that the equalizing bias reflects deliberate journalistic malpractice involving editorial censorship and/or self-censorship [91]. Regardless of which underlying mechanism drives the phenomenon and whether it occurs in an intentional or unconscious manner, the equalizing bias is likely to muddy public understanding of asymmetrical power relations and lead people to suspend judgement and action [28].
@@
Finally, our third finding concerns the presumed reliability of the reported information and the trustworthiness of the source providing it. Unlike AJE, Western journalists were found to engage in strategies aimed at planting seeds of doubt in their readership when reporting on Palestinian victims. When quantifying the human cost of war for civilian victims, Western journalists tended to cite sources more frequently for Palestinian than for Israeli victims, as if Israeli (but not Palestinian) statistics can be taken for granted. This interpretation is further corroborated by a systematic and asymmetrical “casting doubt” style that questions the veridicality of data predominantly when referring to Palestinian victims (such as “reportedly”, “claims”, “according to the Hamas-run health ministry”, and the like). By discrediting sources in a subtle way, journalists undermine the reader’s trust in the information provided and in the reputation of the source.
@@
Together, the present media analysis identifies multiple biases in Western media that may have contributed to the global divide in public opinion and to the delayed public outrage in the Western world in response to the unfolding genocide. Somewhat surprisingly, we did not find a complementary pro-Palestinian bias in AJE, which—with a few exceptions, such as the underreporting of Israeli children’s CVNs—exhibited a relatively balanced reporting pattern.
Scope and Generalizability
Beyond the present case, the study contributes a theory-driven analytic framework for examining how media bias manifests in the coverage of asymmetrical warfare. This framework can be extended to compare media outlets, conflicts (e.g., Israel–Gaza vs. Serbia–Bosnia), or temporal phases within the same conflict (e.g., Gaza wars between 2008 and 2023). To exemplify the latter, we conducted a follow-up replication for our main bias measures to analyze how the same media (AJE, BBC, CNN, NYT) reported on the previous 3 wars on Gaza; for a detailed analysis see Supplementary Note 7. Just like the most recent (2023) war, these three wars, which took place in 2008, 2012, and 2014, showed a strong asymmetry in power and victimhood with Palestinians representing over 90% of all casualties. Importantly, during the 2008 and 2012 wars, Israel was accused of war crimes, but not genocide. In 2014, some organizations voiced genocide accusations, but these did not lead to formal court rulings, unlike the recent 2023 war, where such accusations were brought before the International Court of Justice.
If accusations of ingroup wrong-doing increase the motivation to restore the lost moral ground through competitive victimhood (Noor et al., 2012 [92]; Sullivan et al., 2012[2] , Young & Sullivan, 2016[3] ) and if the equalizing bias offers a way to achieve this, then one would expect a greater equalizing bias the more serious the accusation of wrong-doing. A comparison across the four Gaza wars is broadly consistent with this reasoning. Equalization in casualty reporting appears to have been present in earlier wars and to have intensified over successive conflicts, becoming significantly most pronounced in 2023 (see Supplementary Figure 13), with the notable exception of AJE. It is also in line with another time trend observed in our analysis, namely the tendency to individualize Israeli (vs. Palestinian) victims more clearly in the more recent wars (2014 and 2023) where accusations became more serious (see Supplementary Figure 12). While this interpretation is post-hoc, emphasizing the victimhood of the more powerful group (Israel) remains relevant. As Noor et al. [92] argue, competitive victimhood serves several functions, including inflating the perceived threat from the out-group, justifying in-group violence, and mobilizing moral and material support from third parties. For our last analysis concerning doubt-casting language when reporting casualty statistics, both source-doubting (such as “Hamas-run health ministry”) and doubt-casting expressions (such as “reportedly”) were rare before 2023. They saw a steep increase concerning Palestinian casualty figures in 2023, specifically for BBC. Together, this timeline suggests an increase in Western media bias over time, which is most pronounced for the individualizing and the equalizing biases. This additional analysis illustrates that our framework lends itself to multiple comparisons, including historical trends.